Outdoor Inc recorded the following flex budget and actual results during November 2019:

Actual

Flex Budget

Flex Budget Variance

Percentage Variance

SALES

$215,000

$200,000

COGS

$ 90,000

$ 95,000

Gross Profit

Operating Expenses

$89,000

$81,000

EBIT Earnings Before Interest and Taxes

Interest Expense

$3,850

$3,450

Income Tax

$10,500

$10,500

Net Profit

  1. Fill in the missing amounts in the table (Gross Profit, EBIT and Net Profit)
  2. Calculate flex budget variance and percentage variance. Include ‘favourable F’ or ‘unfavourable U’ rating for each
  3. What type of responsibility centre is shown above, for Outdoor Inc?
  4. If management requires investigation of any variances greater than $8000 and exceeding 10%, which items will be investigated?
Answer #1
a and b) Actual Flex Budget Flex Budget Variance (Actual-Flexbudget) Percentage Variance Favourable/Unfavourable)
SALES $215,000 $200,000 $15000 =(15000/200000) *100 = 7.5 Favourable
COGS $ 90,000 $ 95,000 $5000 =(5000)/95000) *100 = 5.3 Favourable
Gross Profit (Sales - COGS) $125000 $105000 $20000 =(20000/105000)*100=19 Favourable
Operating Expenses $89,000 $81,000 $8000 =(8000/81000)*100=9.9 Unfavourable
EBIT Earnings Before Interest and Taxes (Gross Profit - Operating expenses) $36000 $24000 $12000 =(12000/24000)*100 = 50 Favourable
Interest Expense $3,850 $3,450 $400 =(400/3450)*100 = 11.6 Favourable
Income Tax $10,500 $10,500 $0 0
Net Profit (EBIT-Interest-Income tax) $21650 $10050 $11600 =(11600/10050)*100 = 115.4 Favourable
c) Profit centre
d) Gross profit , EBIT, Interest expense, Net profit variances are more than 10 percent